OpenStudy (anonymous):

tan(pi/5) - tan(pi/30) / 1 + tan(4/5)tan(pi/30)

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

Somebody help

OpenStudy (perl):

$$\large \frac{\tan(\pi/5) - \tan(\pi/30)}{ 1 + \tan(\pi/5) \cdot \tan(\pi/30)} = \tan ( \pi / 5 - \pi /30 ) = ~? $$

OpenStudy (anonymous):

Hold on ima plug it into my calculator

OpenStudy (perl):

you dont need calculator, since you want an 'exact answer' unless your calculator is capable of doing this

OpenStudy (perl):

$$\large \frac{\tan(\pi/5) - \tan(\pi/30)}{ 1 + \tan(\pi/5) \cdot \tan(\pi/30)} = \tan ( \pi / 5 - \pi /30 ) = \tan( \pi / 6 ) $$

OpenStudy (anonymous):

Awww...I wanted to solve it myself...

OpenStudy (anonymous):

But how did you get that end result like that 6

OpenStudy (anonymous):

U divided 30 by 5??

OpenStudy (perl):

$$ \large \frac{\pi }{5} - \frac{\pi}{30} = \color{red}{\frac{6}{6}}\cdot \frac{\pi }{5} - \frac{\pi}{30}= \frac{ 6\pi }{30} - \frac{\pi }{30} = \frac{5\pi }{30} = \frac{\pi }{6}$$

OpenStudy (perl):

yes thats correct

OpenStudy (perl):

divide top and bottom by 5

OpenStudy (anonymous):

Ohhhh ok i get it

OpenStudy (anonymous):

I have another question tho...Im srry

OpenStudy (perl):

so whats tan (pi/6) ?

OpenStudy (anonymous):

5pi/30 or pi/5 - pi/30

OpenStudy (anonymous):

Is that what ur asking??

OpenStudy (perl):

tan pi/6 can be evaluated as a number , a radical

OpenStudy (perl):

just like before tan (45) = 1, tan(pi/6) = ?

OpenStudy (anonymous):

sqtroot2(3) ?

OpenStudy (perl):

tan (pi/6) = sin(pi/6) / cos(pi/6) = ( 1/2 ) / ( sqrt(3) / 2) = 1 / sqrt(3)

OpenStudy (anonymous):

I feel dumb

OpenStudy (perl):

$$ \large \tan (\pi/6) = \frac{\sin(\pi/6)}{ \cos(\pi/6)} = \frac{ 1/2 }{ \sqrt{3} / 2} = \frac{1}{\sqrt 3} = \frac{\sqrt 3 }{3} $$

OpenStudy (anonymous):

so the final answer is sqrtroot(3)/3?

OpenStudy (perl):

yes, with rational denominator

OpenStudy (anonymous):

Wanna help again or its past curfew??

OpenStudy (perl):

sure you can post a new question. if I am around i will look at it.

OpenStudy (anonymous):

k